# Binding Affinity Prediction: From Conventional to Machine Learning-Based Approaches

### v1

@@ 4.3 Ranking Power
Ranking power refers to the ability that a model can correctly rank the know ligands by their binding affinity to the given target protein, usually with the provided poses.
In the ideal case, a binding affinity scoring function/predictor should rank these three ligands accordingly, and therefore the evaluation metrics could be one of Spearman’s rank correlation coefficient ($\rho$) or Kendall’s rank correlation coefficient ($\tau$), defined as the following:
@@
where $r_{i}$ and $r^{\prime}_{i}$ are the true and predicted rank of the $i$th member in the test set of size $n$.
Besides these two commonly used statistical correlation coefficients, some other metrics are used to assess the rank power. For instance, CASF adapts the predictive index (PI) proposed by Pearlman and Charifson (2001) to measure rank, which places a higher weight on the complex pairs with significant differences in binding affinities. PI is computed as:
@@
where $T_{i}$ and $P_{i}$ are the target and predicted binding affinity for the $i$th test sample.
Another example is CI (concordance index) (Gönen and Heller, 2005), used for example by Öztürk et al. (2018) and Öztürk et al. (2019) to measure concordance of target and predicted ordering of binding affinities, computed as:
@@
The latest version of CASF provides a test set with 57 target proteins, each with five ligands that binds the target with distinctively different affinities, resulting in 285 protein-ligand complexes. The standard metrics include Spearman’s rank correlation coefficient ($\rho$), Kendall’s rank correlation coefficient ($\tau$), or Predictive Index (PI).
@@ 4.4 Screening Power
Screening power (Guedes et al., 2018) refers to the ability of a model to tell a true ligand binder from decoy molecules for a given protein target. For the latest version of CASF, there are 57 target proteins and 285 ligands. Each protein has five unique positive ligands, and the other 270 ligands are used as decoy molecules. For each protein and each decoy molecule, the authors have collected 100 potential binding poses from various softwares. The tested scoring function or binding affinity predictor will be applied to every pose, and the highest score/affinity will be considered as the score/affinity of the this protein-ligand pair. And finally, the screening power of the model is tested by ranking all the ligands in the descending order of their score/affinity and see if all five true ligands are ranked in the top, which is characterized with enhancement factor (EF), computed as:
@@
where $\alpha$ is an arbitrary percentage threshold, typically 1%, 5%, or 10%; $\textit{NTB}_{\alpha}$ is the number of true binders found among the top $\alpha$ candidates; and $\textit{NTB}_{\text{total}}$ is the total number of true binders for a single target protein, which in CASF’s case is 5. A major problem with CASF’s approach is that they assumed that random ligands in the dataset do not bind with a target protein, which is often not the case. In fact, Su et al. (2018) report that 21 of the 57 target proteins have more than five binding ligands in the dataset according to known binding data in ChEMBL, which makes CASF’s evaluation for screening power rather questionable. Moreover, using random molecules as non-binders in a decoy set is a common practice in the research of protein-ligand binding affinity and drug-target interaction, which could potentially hinder the learning and validation for the models.
@@
Among all four evaluation aspects, scoring power, on the surface level, seems to be the dominant factor over the other three. Essentially, docking power, ranking power, and screening power are evaluated with the scores/affinities predicted from the model. Thus, it is completely reasonable to assume that a binding affinity predictor with good scoring power might also perform well on the other three evaluations. However, it seems that in the literature such as (Cheng et al., 2009), it is simply not the case. For conventional methods, we usually found that ranking is a much harder task and for machine learning based predictors, despite their impressive results for scoring power, they seem to perform poorly on docking, which renders the whole category of methods unconvincing.
5 Conclusion
We have provided an extensive review of binding affinity prediction methodologies, tracing their evolution from conventional approaches to the latest machine learning (ML) and deep learning (DL) techniques. Binding affinity, which refers to the strength of interaction between a small molecule (like a drug) and a target protein, is crucial in drug discovery and other areas of bioinformatics. Our review has highlighted the increasing application of ML and DL methods in this domain, driven by the growing availability of protein and small molecule data.
We categorized binding affinity prediction methods into three main types: conventional methods, traditional machine-learning-based methods, and deep-learning-based methods. Conventional methods rely heavily on established chemical knowledge and physics-based calculations, often resulting in rigid models that perform well only in specific cases. In contrast, traditional ML methods improve prediction accuracy by leveraging human-engineered features, while ML methods, which have gained prominence more recently, aim to automate feature extraction and can potentially outperform other methods as more data becomes available.
We also discussed challenges in binding affinity prediction, such as the limitations of available datasets, which often contain biased or insufficient data, and the inherent complexity of protein-ligand interactions that make accurate predictions difficult. We emphasize the need for more robust datasets and models that can handle the dynamic nature of protein-ligand interactions.
@@ 5 Conclusion
Our review also leads us to identify five important areas for further research, as follows. 1) Overcome limitations in existing datasets. Current datasets are often biased towards protein-ligand pairs with favorable binding constants and correct poses, while complexes with low affinity or failed bindings are underrepresented. Future research could focus on creating more balanced datasets that better represent the full spectrum of binding affinities. 2) Integrate physics-based models with ML/DL. There is potential for hybrid models that combine the theoretical rigor of physics-based methods with the predictive power of machine learning. Research in this area could explore how best to integrate these approaches to improve prediction accuracy. 3) Better handle protein flexibility. Most existing models treat proteins as rigid entities, which is a significant simplification. Developing methods that account for protein flexibility, as well as different conformation of ligand, during binding could lead to more accurate predictions. 4) Improve evaluation metrics. Current evaluation metrics, such as Pearson correlation for scoring power and RMSD for docking power, have limitations. Research could explore alternative metrics that better capture the nuances of binding affinity predictions, particularly in real-world applications like drug discovery. 5) Multimodal data integration. With the growing interest in multimodal models, future research could focus on integrating diverse data types (e.g., sequence data, 3D structures, interaction networks) to capture complex interactions more effectively. Research in each of these areas promises to advance the field of binding affinity prediction, potentially leading to more accurate and reliable tools for drug discovery and other applications.

### v2

@@ 4.3 Ranking Power
Ranking power refers to a model’s ability to correctly rank the ligands based on their binding affinity to the given target protein. A binding affinity scoring function/predictor may rank ligands with Spearman’s rank correlation coefficient ($\rho$) or Kendall’s rank correlation coefficient ($\tau$) as the evaluation metrics:
@@
where $r_{i}$ and $r^{\prime}_{i}$ are the true and predicted rank of the $i$th member in the test set of size $n$. Besides these two commonly used statistical correlation coefficients, some other metrics are used to assess the rank power. For instance, CASF adapts the predictive index (PI) proposed by Pearlman and Charifson (2001) to measure rank, which places a higher weight on the complex pairs with significant differences in binding affinities:
@@
where $T_{i}$ and $P_{i}$ are the target and predicted binding affinity for the $i$th test sample. Another example is CI (concordance index) (Gönen and Heller, 2005), used to measure concordance of target and predicted ordering of binding affinities (Öztürk et al., 2018, 2019):
@@ 4.4 Screening Power
Screening power (Guedes et al., 2018) refers to a model’s ability to differentiate the true ligand binder from decoy molecules for a given protein target. Screening power is computed by ranking all of the ligands in descending order of their score/affinity, and determining whether the true ligands are ranked highly, which is characterized with enhancement factor (EF):
@@
where $\alpha$ is an arbitrary percentage threshold, typically 1%, 5%, or 10%; $\textit{NTB}_{\alpha}$ is the number of true binders found among the top $\alpha$ candidates; and $\textit{NTB}_{\text{total}}$ is the total number of true binders for a single target protein. A major problem is the assumption that random ligands in the dataset do not bind with a target protein, which is often not the case. For instance, Su et al. (2018) report that 21 of the 57 target proteins ranked by CASF have more than five binding ligands in the dataset, according to known binding data in ChEMBL, which raises questions regarding CASF’s evaluation for screening power. Moreover, using random molecules as non-binders in a decoy set—although considered common practice in protein-ligand binding affinity and drug-target interaction modeling—can potentially hinder model learning and validation.
@@
Among the four evaluation aspects, scoring power seems to be the dominant factor, at least on the surface level. Essentially, docking power, ranking power, and screening power are evaluated with the scores/affinities predicted from the model. Thus, it is completely reasonable to assume that a binding affinity predictor with good scoring power might also perform well on the other three evaluations. However, it may not be the case (Cheng et al., 2009). Ranking is a much harder task for conventional methods, and machine learning based predictors—despite their impressive results for scoring power—seem to perform poorly on docking, which renders the whole category of methods unconvincing.
5 Discussion
We have provided an extensive review of binding affinity prediction methodologies, tracing their evolution from conventional approaches, traditional machine-learning-based methods, to modern deep-learning-based methods. We also examine key challenges in binding affinity prediction, including dataset limitations such as measurement inconsistencies, chemical space biases, and incomplete experimental coverage, as well as the intrinsic complexity of protein-ligand interactions that collectively impede accurate computational prediction.
@@ 5 Discussion
Our review also leads us to identify five important areas for further research. 1) Overcome limitations in existing datasets. Current datasets are often biased towards protein-ligand pairs with favorable binding constants and correct poses, while complexes with low affinity or failed bindings are underrepresented. Future research could focus on creating more balanced datasets that better represent the full spectrum of binding affinities. 2) Integrate physics-based models with machine learning and deep learning. There is potential for hybrid models that combine the theoretical rigor of physics-based methods with the predictive power of machine learning. Research in this area could explore how to better integrate these approaches to improve prediction accuracy. 3) Better handle protein flexibility. Most existing models treat proteins as rigid entities, a simplification that overlooks important dynamics. Developing methods that account for protein flexibility, along with different ligand conformations during binding, could yield more accurate predictions. 4) Improve evaluation metrics. Current evaluation metrics, such as Pearson correlation for scoring power and RMSD for docking power, have limitations. Research could explore alternative metrics that better capture the nuances of binding affinity predictions, particularly in real-world applications like drug discovery. 5) Multimodal data integration. With the growing interest in multimodal models, future research could focus on integrating diverse data types (e.g., sequence data, 3D structures, interaction networks, molecular dynamics) to capture complex interactions more effectively. Research in each of these areas can further advance the field of binding affinity prediction, potentially leading to more accurate and reliable tools for drug discovery and other applications.
Further, with the Food and Drug Administration (FDA) recently phasing out animal testing (Williams, 2024), AI-driven tools—especially virtual cells (AIVCs) (Bunne et al., 2024; Song et al., 2024)—offer a transformative, multiscale approach to simulating and analyzing molecules, cells, and tissues. The development of AIVCs offers a promising path forward by embedding molecular binding in a dynamic, data-integrated, and biologically realistic context. This systems-level perspective advances binding affinity prediction beyond isolated structural modeling, improving both its accuracy and translational relevance in computational drug discovery. We believe AIVCs can be leveraged to enhance binding affinity prediction in several key ways. 1) Contextual environment simulation. Traditional models often assess binding in simplified, isolated systems. In contrast, AIVCs simulate full cellular environments, enabling context-aware predictions for complex targets, such as membrane proteins, and allosteric sites. 2) Dynamic protein conformations. Binding is inherently dynamic. AIVCs can capture ligand entry and exit pathways, conformational transitions, and temporal fluctuations, providing a more realistic representation of binding processes. 3) Multi-target and off-target analysis. By representing the entire proteome and interactome, AIVCs can evaluate binding specificity and off-target risks in parallel, improving early-stage selectivity assessments. 4) Integration of multi-omics and systems constraints. AIVCs can incorporate transcriptomic, proteomic, metabolomic, and genomic data—enabling simulation of protein abundance, isoform specificity, metabolic competition, and mutation effects. This facilitates personalized affinity predictions and system-aware drug efficacy assessments by enabling cell type—or patient-specific predictions—key to advancing personalized medicine and precision pharmacology.
Enabled by innovative dataset curation, novel algorithmic development, and creative evaluation, we envision an emerging paradigm shift in binding affinity prediction to advance precision pharmacology.
Acknowledgements
We thank Nabil Faruk for constructive suggestions. This work is supported in part by the RadBio-AI project (DE-AC02-06CH11357), U.S. Department of Energy Office of Science, Office of Biological and Environment Research, the Improve project under contract (75N91019F00134, 75N91019D00024, 89233218CNA000001, DE-AC02-06-CH11357, DE-AC52-07NA27344, DE-AC05-00OR22725), the Exascale Computing Project (17-SC-20-SC), a collaborative effort of the U.S. Department of Energy Office of Science and the National Nuclear Security Administration.